# Optimizing Distributed Reinforcement Learning with Reactor Model and Lingua Franca

### 07 Dec 2023 v1

@@ 3.3 Ray
Ray liang2021rllib ; moritz2018ray is an open-source framework designed for distributed computing and provides a simple, universal API for building distributed applications. Ray utilizes the Actor model, which is central to its architecture and operation, making it highly effective for concurrent and distributed computing tasks. The Actor model in Ray is used to encapsulate state and behavior, with actors being distributed across a node or a cluster and communicating through asynchronous message passing. Actors can be used for performance reasons (like caching soft state or ML models), or they could be used for managing long living connections to databases or to web sockets. They can maintain state across multiple tasks, which is particularly useful for applications that require managing large, mutable states, such as machine learning models or large datasets.
@@
The local scheduler in Ray is a component of a worker node called Raylet. Raylet manages the worker processes and consists of two components, a task scheduler and an object store. The task scheduler takes care of scheduling and executing work on a node. It addresses issues such as a worker being busy, not having the proper resources to run a task, or not having the values it needs to run a given task. Expensive serialization and deserialization as well as data copying are a common performance bottleneck in distributed computing. Shared memory, specifically that which is managed directly by the operating system kernel, emerges as a superior mechanism compared to conventional approaches like socket connections. This kind of memory management boasts efficiency that transcends even memory-mapped file-based sharing. The fundamental attribute of distributed object store is that all objects within the store are immutable and retained in shared memory. The object store has its eviction policy, removing objects from the store or transferring them to other nodes when the allocated size limit is exceeded. This design choice ensures optimal access speeds, particularly when multiple workers on a singular node need to engage with the data. It’s noteworthy that every node possesses an independent object store. Each node of a Ray is equipped with an object store, within that node’s Raylet, and all objects stored collectively form the distributed object store of a cluster. The object store manages a shared pool of memory across workers on the same node. The object store is implemented in Plasma, which belongs to the Apache Arrow project. Functionally, the object store takes care of memory management and ultimately makes sure workers have access to the objects they need in Ray.
3.4 Lingua Franca
@@
Lingua Franca (LF) menard2023high ; schulz2023polyglot is a polyglot coordination language designed to facilitate the development of concurrent systems by focusing on deterministic interactions with their environment. It operates on the reactor model, where reactors are the fundamental units of composition, each encapsulating reactions to external stimuli. LF’s main feature is its deterministic nature, meaning that given a set of inputs, a LF program will always produce the same outputs, greatly enhancing testability and efficiency. This determinism is achieved by using a superdense model of time (where events are ordered by time and microstep), ensuring causality and the absence of non-deterministic feedback loops in the reaction network.
@@
The architecture of LF includes a compiler (lfc). The compiler process involves parsing and validating the LF code, checking for syntax errors, instantiation cycles, and cyclic dependencies among reactions. Valid code can then be transpiled into target code in languages like C, C++, TypeScript, and Python, which is then combined with a runtime system to manage the execution of reactors. In this paper, we focus on the Python target as most RL applications are written in Python. LF also allows for graphical program representation, enhancing program structure understanding and error identification.
@@
LF allows for flexibility in connecting reactors through multiports (ports handling multiple data channels) and banks (multiple reactor instances). Connections can be logical, implying synchronization between ports, or physical, introducing intentional non-determinism for scenarios where strict event ordering is not required. LF’s syntax permits the creation of complex interaction patterns among reactors, offering a robust toolset for constructing scalable and maintainable concurrent systems.
The formal semantics of LF are grounded in the theory of discrete-event systems, utilizing a generalized ultrametric space for modeling the behavior of LF programs. This approach guarantees that each LF program is deterministic and adheres to causality, essential for reliability in concurrent system design. LF’s semantic model is fully abstract, providing both an operational perspective (how the program executes) and a denotational perspective (the meaning of program constructs), ensuring consistency.
3.5 Representing Distributed RL As Dataflow Graph with Lingua Franca
Figure 1: Generated Dataflow Graph for distributed RL
LF provides a streamlined process to represent distributed RL as dataflow graphs lohstroh2021toward . By simply initializing reactors and setting their input and output ports, LF can automatically generate the corresponding dataflow graph. This diagrammatic feature is seamlessly integrated with Visual Studio extensions and Eclipse, offering a visual representation of the underlying RL processes. The resulting diagram (Fig. 1) clearly displays the reactors, ports, state variables, and their data dependencies, simplifying the understanding of the system’s structure and aiding in the debugging of distributed RL systems. Users are therefore better positioned to manage the parallelization of RL processes. Detailed information about the implementation is provided in the appendix.
@@ 3.5 Representing Distributed RL As Dataflow Graph with Lingua Franca
  • PolicyState: The internal state of the policy, significant for stateful policies such as those using recurrent neural networks.
@@
ReplayBufferReactor: Acting as a centralized experience replay buffer, this reactor stores trajectories from the RolloutReactor for subsequent sampling. It provides sampled experience batches to the LearnerReactor for policy updates.
@@
  • SamplingPointer: Indices or pointers facilitating efficient experience sampling.
  • PrioritizedInfo: For prioritized replay buffers, additional data to manage experience priorities.
LearnerReactor: The LearnerReactor updates policies based on sampled experiences from the ReplayBufferReactor or directly from the RolloutReactor, and broadcast updated gradients to the RolloutReactor.
@@
  • OptimizerState: Elements related to the optimization process, such as momentum and learning rates.
@@
  • TargetNetworkParameters: In algorithms like DQN mnih2013playing , these are slowly updated parameters offering a stable learning target.
4 Optimizations for Distributed Reinforcement Learning
4.1 Scheduling Algorithm
@@
The execution process of LF programs involves a scheduler responsible for overseeing all scheduled future events, managing the logical time progression, and executing triggered reactions in the order dictated by the dependency graph. The scheduling mechanism in the LF model, depicted in fig. 2, operates with an event queue that strictly follows a tag order for processing upcoming events. Upon adding an event to the queue and processing it, the scheduler identifies triggered reactions, placing them in a reaction queue. These reactions are subsequently moved to a ready queue and executed by worker threads once all dependencies, as outlined by the Action-Port Graph (APG), are satisfied. Within this scheduling framework, the scheduler ensures a sequential execution of reactions, considering dependencies and aiming to maximize parallelism.
@@
The scheduling mechanism is similar to Directed Acyclic Graph (DAG)-based strategies but differs by accommodating reactions within the Action-Port Graph (APG) that may not always need execution. Since the scheduler cannot predict in advance which reactions will be triggered at a given tag, it cannot precompute an optimal schedule. To address this, the scheduler assigns a level to each reaction, allowing reactions at the same level to execute concurrently. This heuristic eliminates the need for runtime APG analysis, streamlining the process. The scheduler processes reactions in a stepwise way, advancing to the next level only after completing all reactions at the current level. This approach substantially reduces synchronization overhead and contention on shared resources, contributing to its efficiency.
Figure 2: Scheduling mechanism in the LF runtime  menard2023high
4.2 Scheduling Optimizations
@@
The scheduling algorithm described in the above section is relatively straightforward to implement, but achieving competitive performance requires additional optimizations. These optimizations, implemented in the C runtime, are detailed below.
@@
Coordination of worker threads: Conceptually, the scheduler and workers are distinct; however, in practice, having a central scheduler with separate worker threads can cause significant synchronization overhead. To mitigate this, our implementation allows any worker thread to become the scheduler, allowing it to move ready reactions to the queue or advance to the logical time once all reactions are processed. Additionally, LF leverage the fact that we know the number of parallel reactions to execute from the APG, and thus can use a counting semaphore to regulate the number of active workers.
@@
Lock-free data structure: The three queues (event, reaction, and ready queue) and other data structures are shared across all workers. Mutex-based synchronization would be inefficient due to high contention over these shared resources. Accordingly, LF adopted lock-free data structures wherever feasible. For example, the ready queue is a fixed-size buffer with an atomic counter, which corresponds to the maximum number of parallel reactions defined by the APG level. Each time a worker tries to execute a reaction it atomically decrements the counter. If the counter is negative, it indicates an empty queue, and the worker should proceed accordingly. Otherwise, the counter provides the index within the buffer from which to read. This operation is safe without additional synchronization, as all workers are waiting for upcoming new reactions.
4.3 Enabling Multi-threading for Reinforcement Learning
@@
  3. 3.
    I/O Bound Task Optimization: Multithreading proves advantageous in I/O-bound operations, where the ability to perform other tasks while waiting for I/O operations enhances efficiency.
  4. 4.
@@
    Multi-Core Utilization: Multiprocessing enables the full utilization of multi-core processors by running separate processes on each core. However, it’s important to consider that when sending excessively large objects, the processor may spend a significant amount of time on I/O overhead. LF, on the other hand, is able to bring threads to its full utilization
4.4 Optimizing Thread Allocations
@@
The principle that running a task with multiple threads can speed up the process is generally true, as each thread can handle a portion of the work simultaneously. However, the scaling is not always linear due to various factors like thread allocation and CPU architecture. In our benchmarks, the Intel CPU has a hybrid architecture, featuring a combination of performance and efficiency cores. Performance cores are designed for high-speed and intensive tasks, while efficiency cores are optimized for lower power consumption and handling background tasks. Current implementation of LF runtime randomly allocates threads for workers. As a result, when synchronous tasks are distributed across these cores, the overall speed is bounded by the slower efficiency cores. To achieve linear scaling, the Reactor C runtime has been optimized to prioritize using the performance cores before spawning threads.
5 Performance Comparison Between LF and Ray
The actor model is a popular approach for developing distributed RL applications, with Ray being notable for their usability and efficiency in handling multiple actors. However, Lingua Franca (LF) introduces a different model of computation, imposing more restrictions compared to the actor model. We hypothesize that the reactor model can outperform the traditional actor models used in distributed RL frameworks (e.g. Ray). Drawing evidence from this study, the reactor model has been shown to surpass the performance of the traditional Actor Framework, Akka and C++ Actor Framework, by factors of 1.86x and 1.42x, respectively. More specifically, the reactor model with a fixed set of actors and fixed communication patterns allows the scheduler to eliminate works needed for synchronization (e.g. acquiring and releasing locks for each actor, or sending and processing coordination-related messages between environment and policy). Furthermore, LF leverages No GIL Python for multithreading, which offers several benefits, including efficient data sharing within a shared memory space, resource efficiency due to lower resource consumption compared to processes, and faster context switching thanks to the shared process and memory space. The reduced overhead is demonstrated by both figure 7 and 8. All measurements were performed on AWS EC2 m5.8xlarge instance, equipped with an Intel® Xeon® Platinum 8175M CPU @ 2.50GHz featuring 32 vCPUs. This setup includes 128 GiB of RAM and offers a 10 Gbps network bandwidth. The system runs on Ubuntu 20.04 and use Python 3.9.10, with numpy version 1.22.3 and gym version 0.19.0. Benchmarks are available at this GitHub repository.
@@ 5 Performance Comparison Between LF and Ray
Figure 3: Mean Overhead of Broadcast and Gather 10MB Object with Different Number of Actors using Ray and LF.
Figure 3: Mean Overhead of Broadcast and Gather 10MB Object with Different Number of Actors using Ray and LF.
@@
Figure 4: Mean Overhead of Broadcast and Gather on 16 actors with Different Object Sizes using Ray and LF.
5.1 Number of Actors

### 02 Feb 2024 v2

@@ 3.3. Ray
Ray (Liang et al., 2021; Moritz et al., 2018) is an open-source framework designed for distributed computing and has been widely used to parallelize RL workloads from single-node to multi-nodes. Ray utilizes the actor model, which is central to its architecture and operation, making it highly effective for concurrent and distributed computing tasks. The actor model in Ray is used to encapsulate state and behavior, with actors being distributed across a node or a cluster and communicating through asynchronous message passing. It is important to highlight that in this study, our primary focus is on enhancing the efficiency of distributing actors within a node. Actors can be used for performance reasons (like caching soft state or ML models), or they could be used for managing long-living connections to databases or to web sockets. They can maintain state across multiple tasks, which is particularly useful for applications that require managing large, mutable states, such as machine learning models or large datasets.
@@
The local scheduler in Ray is a component of a worker node called Raylet. Raylet manages the worker processes and consists of two components, a task scheduler and an object store. The task scheduler takes care of scheduling and executing work on a node. It addresses issues such as a worker being busy, not having the proper resources to run a task, or not having the values it needs to run a given task. Expensive serialization and deserialization as well as data copying are common performance bottleneck. Shared memory, specifically that which is managed directly by the operating system kernel, emerges as a superior mechanism compared to conventional approaches like socket connections. The fundamental attribute of the in-memory object store is that all objects within the store are immutable and retained in shared memory. The object store has its eviction policy, removing objects from the store or transferring them to other nodes when the allocated size limit is exceeded. This design choice ensures optimal access speeds, particularly when multiple workers on a singular node need to engage with the data. Each node provisioned by Ray is equipped with an object store, within that node’s Raylet. Functionally, the object store takes care of memory management and ultimately makes sure workers have access to the objects they need in Ray.
3.4. Lingua Franca
@@
Lingua Franca (LF) (Menard et al., 2023; Schulz-Rosengarten et al., 2023) is a polyglot coordination language designed to facilitate the development of concurrent systems by focusing on deterministic interactions with the environments. It operates on the reactor model, where reactors are the fundamental units of composition, each encapsulating reactions to external stimuli. LF’s main feature is its deterministic nature, meaning that given a set of inputs, a LF program will always produce the same outputs, greatly enhancing testability and efficiency. This determinism is achieved by using a superdense model of time (where events are ordered by time and microstep), ensuring causality and the absence of non-deterministic feedback loops in the reaction network.
@@
The architecture of LF includes a compiler (lfc). The compiler process involves parsing and validating the LF code, checking for syntax errors, instantiation cycles, and cyclic dependencies among reactions. Valid code can then be transpiled into target code in languages like C, C++, TypeScript, and Python, which is then combined with a runtime system to manage the execution of reactors. LF also allows for graphical program representation, enhancing program structure understanding and error identification.
@@
LF allows for flexibility in connecting reactors through multiports (ports handling multiple data channels) and banks (multiple reactor instances). Connections can be logical, implying synchronization between ports, or physical, introducing intentional nondeterminism for scenarios where strict event ordering is not required. LF’s syntax permits the creation of complex interaction patterns among reactors, offering a robust toolset for constructing scalable and maintainable concurrent systems.
The formal semantics of LF is grounded in the theory of discrete-event systems, utilizing a generalized ultrametric space for modeling the behavior of LF programs (Liu et al., 2006). This approach guarantees that each LF program is deterministic and adheres to causality, essential for reliability in concurrent system design. LF’s semantic model is fully abstract, providing both an operational perspective (how the program executes) and a denotational perspective (the meaning of program constructs), ensuring consistency. An operational semantics of the reactor model based on a formalization in LEAN (de Moura et al., 2015) is also available (Rossel et al., 2023).
Figure 1. LF: compilation process
Fig.  1 depicts the compilation process of the LF framework. This process begins with the source code written in Lingua Franca, with a .lf extension. The source code is then processed by a modified lfc compiler, tailored to work without Python’s Global Interpreter Lock (GIL). Additionally, the compilation process involves various no-GIL versions of the Python libraries such as NumPy, PyTorch, Gym, etc.
This compiler generates C code for each reactor, with files named R1.c, R2.c, and so on as well as the compiled Lingua Franca library, LinguaFranca.so, which can be imported and used in Python scripts (*.py). The runtime implementation serves as a bridge between the high-level coordination language and the underlying Python environment without the GIL, enabling users to write truly concurrent Python programs. It abstracts away the burden of coordinating worker threads.
3.5. Representing RL tasks as a Dataflow Graph
The diagram synthesizer in LF provides a streamlined process to represent parallel RL tasks as dataflow graphs (Lohstroh et al., 2021). By simply initializing reactors and setting their input and output ports, LF can automatically generate the corresponding dataflow graph. This diagrammatic feature is seamlessly integrated with Visual Studio Code and Eclipse, offering a visual representation of the underlying RL processes. The resulting diagram (Fig. 2) clearly displays the reactors, ports, state variables, and their data dependencies, simplifying the understanding of the system’s structure and aiding in the debugging of parallel RL tasks. Users are therefore better positioned to manage the parallelization of RL processes. LF also offers a compact syntax for ports that are capable of sending or receiving across various channels, and a syntax for multiple instances of a reactor. These concepts are known as multiports and banks of reactors. For example, we have created a bank of six instances of ReplayBufferReactor and one instance of LearnerReactor, connecting them using multiport. Detailed information about the implementation is provided in the appendix.
@@ 3.5. Representing RL tasks as a Dataflow Graph
  • PolicyState: The weight of the current policy.
@@
ReplayBufferReactor: Acting as a centralized experience replay buffer, this reactor stores trajectories from the RolloutReactor for subsequent sampling. It provides batched experience to the LearnerReactor for policy updates.
@@
  • SamplingPointer: Indices or pointers facilitating efficient sampling.
  • PrioritizedInfo: For prioritized replay buffers, additional information will be used to manage sampling.
LearnerReactor: The LearnerReactor updates policies based on sampled experiences from the ReplayBufferReactor or directly from the RolloutReactor, and broadcasts updated gradients to the RolloutReactor.
@@
  • OptimizerState: Elements related to the optimization process, such as momentum variables and correction terms.
@@
  • TargetNetworkParameters: In algorithms like DQN (Mnih et al., 2013), these are slowly updated parameters offering a stable learning target.
Figure 2. Generated Dataflow Graph for parallel RL tasks
4. Optimizations for Parallel Reinforcement Learning
4.1. Scheduling Algorithm
@@
The execution process of programs in LF involves a scheduler responsible for overseeing all scheduled future events, managing the logical time progression, and executing triggered reactions in the order dictated by the dependency graph. The scheduling mechanism in LF, depicted in Fig. 3, operates with an event queue that strictly follows a tag order for processing upcoming events. Upon adding an event to the queue and processing it, the scheduler identifies triggered reactions, placing them in a reaction queue. These reactions are subsequently moved to a ready queue and executed by worker threads once all dependencies, as outlined by the Action-Port Graph (APG), are satisfied. Within this scheduling framework, the scheduler ensures a sequential execution of reactions that depend on one another, aiming to maximize parallelism.
@@
The scheduling mechanism is similar to Directed Acyclic Graph (DAG)-based strategies but differs by accommodating reactions within the Action-Port Graph (APG) that may not always need execution. Since the scheduler cannot predict in advance which reactions will be triggered at a given tag, it cannot precompute an optimal schedule. To address this, the scheduler assigns a level to each reaction, allowing reactions at the same level to execute concurrently. This heuristic eliminates the need for runtime APG analysis, streamlining the process. The scheduler processes reactions sequentially, advancing to the next level only after completing all reactions at the current level. This approach substantially reduces synchronization overhead and contention on shared resources, contributing to its efficiency.
Figure 3. Scheduling mechanism in the LF runtime  (Menard et al., 2023)
4.2. Scheduling Optimizations
@@
The scheduling algorithm described in the above section is relatively straightforward to implement, but achieving competitive performance requires additional optimizations. These optimizations, implemented in the C runtime in LF, are detailed below.
@@
Coordination of worker threads: Conceptually, the scheduler and workers are distinct; however, in practice, having a central scheduler with separate worker threads can cause significant synchronization overhead. To mitigate this, our implementation allows any worker thread to become the scheduler, allowing it to move ready reactions to the queue or advance to the logical time once all reactions are processed. Additionally, LF leverages the fact that we know the number of parallel reactions to execute from the APG, and thus can use a counting semaphore to regulate the number of active workers.
@@
Lock-free data structure: The three queues (event, reaction, and ready queue) and other data structures are shared across all workers. Mutex-based synchronization would be inefficient due to high contention over these shared resources. Accordingly, LF adopts lock-free data structures wherever feasible. For example, the ready queue is a fixed-size buffer with an atomic counter, which corresponds to the maximum number of parallel reactions defined by the APG level. Each time a worker tries to execute a reaction it atomically decrements the counter. If the counter is negative, it indicates an empty queue, and the worker should proceed accordingly. Otherwise, the counter provides the index within the buffer from which to read. This operation is safe without additional synchronization, as all workers are waiting for upcoming new reactions.
4.3. Enabling Multi-threading for Reinforcement Learning
@@
  3. (3)
    I/O Bound Task Optimization: Multithreading proves advantageous for I/O-bound operations, where the ability to perform other tasks while waiting for I/O operations enhances efficiency.
  4. (4)
@@
    Multi-Core Utilization: Multiprocessing enables efficient utilization of multi-core processors by running separate processes on each core. However, it’s important to consider that when sending excessively large objects, the processor may spend a significant amount of time on I/O overhead.
4.4. Optimizing Thread Allocations
@@
The principle that running a task with multiple threads can speed up the process is generally true, as each thread can handle a portion of the work simultaneously. However, the scaling is not always linear due to various factors like thread allocation and CPU architecture. In our benchmarks, the Intel CPU has a hybrid architecture, featuring a combination of performance and efficiency cores. Performance cores are designed for high-speed and intensive tasks, while efficiency cores are optimized for lower power consumption and handling background tasks. The original LF runtime randomly allocates threads for workers. As a result, when synchronous tasks are distributed across these cores, the overall speed is bounded by the slower efficiency cores. To achieve linear scaling, the Reactor C runtime has been optimized to prioritize using the performance cores before spawning threads.
@@
Figure 4. Mean Overhead of Broadcast and Gather 10MB Object with Different Number of Actors using Ray and LF.
@@
Figure 5. Mean Overhead of Broadcast and Gather on 16 actors with Different Object Sizes using Ray and LF.
5. Performance Comparison Between LF and Ray
To demonstrate the improved performance of LF over Ray for parallel RL workloads, we run the following experiments. We believe these experiments are comprehensive and adequate to support our assertion that our optimized implementation of the reactor model for RL significantly advances the state of the art for parallel RL.
  1. (1)
    We demonstrate lower overhead of broadcast and gather operations with LF than with Ray, and show that the overhead difference scales favorably for LF as we increase both the number of actors/reactors and the communication object size. While not an RL benchmark, this toy workload resembles the computation done in RL, and is useful for highlighting why exactly it is that LF outperforms Ray (sections 5.2, 5.3).
  2. (2)
    In most parallel RL settings, (e.g., as is typically the case with policy gradient and Q-learning algorithms), the majority of the parallelism that can be obtained by porting these algorithms to leverage parallel compute is in the data collection phase of the algorithm (as opposed to for example value estimation or policy improvement). As such, we extensively test parallel data collection in popular RL simulated environments with both LF and Ray (sections 5.4, 5.5).
  3. (3)
    To demonstrate that LF integrates seamlessly with other forms of parallel compute paradigms, namely GPU acceleration, and to demonstrate the versatility LF provides to implement various reinforcement learning algorithms, we implement and evaluate synchronized parallel Q-learning with deep neural networks (section 5.6).
  4. (4)
    Finally, we observe that multi-agent RL lends itself very well to being parallelized, as each agent can maintain its own learning and data collection actors/reactors. We evaluate in such a multi-agent RL (MARL) setting and observe favorable results for LF (section 5.7).
5.1. Experimental Setup
The actor model is a popular approach for developing parallel RL applications, with Ray being notable for its usability and efficiency in handling multiple actors. However, LF introduces a different model of computation, imposing more restrictions compared to the actor model. We hypothesize that the reactor model with LF can outperform Ray, which is very widely used today for parallel RL. Drawing evidence from this study (Menard et al., 2023), the reactor model has been shown to surpass the performance of the traditional Actor Framework, Akka and C++ Actor Framework, by factors of 1.86x and 1.42x, respectively, for non-RL workloads. More specifically, the reactor model with a fixed set of actors and fixed communication patterns allows the scheduler to eliminate work needed for synchronization. Furthermore, with our customized implementation, LF leverages No GIL Python for multithreading, which offers several benefits, including efficient data sharing within a shared memory space and faster context switching. The reduced overhead is demonstrated by both Fig. 4 and Fig. 5 . The term, overhead, refers to the duration required for one actor to send a payload to another and for it to be received. This process may include serialization and deserialization, acquiring and releasing locks for each actor, sending and processing coordination-related messages, as well as transferring data over the network. All measurements were performed on AWS EC2 m5.8xlarge instance, equipped with an Intel® Xeon® Platinum 8175M CPU @ 2.50GHz featuring 32 vCPUs. This setup includes 128 GiB of RAM and offers a 10 Gbps network bandwidth. The system runs on Ubuntu 20.04 and use Python 3.9.10, with NumPy version 1.22.3 and gym version 0.19.0.
5.2. Number of Actors